A 500g sample of naphthalene (MW=128.17g/mol, kf=6.9°C/m, tpure=80.6°C) also contains an additional 109g of an unknown chemical (i=1). The freezing point of this solution was 65.8°C. What is the molar mass of this unknown substance?

Please help! I have no idea how to do this.
Answer #1

Actual freezing point of Naphthalene = 80.6 °C

The following formula can be used

Δt = i Kf m

m = molality of the solution = X / 0.5 kg

X = moles of

Kf for Naphthalene is 6.9 °C/m

i = 1

Δt = 80.6 °C - 65.8°C =  14.8°C

Substitute the values in the equation

14.8°C = 1 x 6.9 °C/m x (x / 0.5 kg)

x = 1.072 Moles

Molar mass of the compound = 109 g / 1.072 moles = 101.635 g/mol